# ViSA-Flow: Accelerating Robot Skill Learning via Large-Scale Video Semantic Action Flow

### v2

@@ Abstract
11footnotetext: Equal contribution. Corresponce to: changhec@umich.edu, quantao@kth.se.
Keywords: Robot Manipulation, Imitation Learning, Learning from Video
1 Introduction
@@
Robot imitation learning has achieved remarkable success in enabling robots to acquire complex manipulation skills, ranging from basic object manipulation [1, 2] to intricate assembly procedures [3]. However, the scalability of traditional imitation learning approaches is fundamentally limited by the need for extensive, carefully curated robot datasets that are costly to collect. This has become a critical bottleneck in developing robots capable of performing diverse real-world tasks.
@@
In contrast, humans demonstrate an extraordinary ability to learn new skills by observing others. Whether it be in person, instructional videos or even from sports broadcasts, humans instinctively focus on the semantically relevant components. For instance, when learning tennis, we naturally attend to the player’s body movements, racquet handling techniques, and ball trajectories, while effectively filtering out irrelevant background information. This selective attention to meaningful elements enables efficient skill acquisition and transfer. The vast repository of publicly available videos on the internet similarly represents an untapped resource for robot learning, offering diverse demonstrations of human skills across countless domains. However, effectively leveraging this resource requires addressing several key challenges, particularly in bridging the gap between human demonstrations in unconstrained videos and robot execution in the real world.
@@
Figure 1: Learning Robot Manipulation Skills from Human Videos via Semantic Action Transfer. (a) Humans and robots often share underlying atomic actions for similar tasks (e.g., Move & Grasp). (b) Our framework leverages large-scale, unlabeled human videos by extracting weakly supervised semantic action flow priors (ViSA-Flow). This knowledge is distilled into a human policy and efficiently transferred to learn a corresponding robot policy.
@@
Recent research [4, 5, 6] has explored enabling robots to acquire skills by directly observing unstructured human videos. These approaches have demonstrated strong generalizability, allowing robots to adapt to new tasks effectively. In most real-world scenarios, when humans learn a skill, we primarily focus on the interaction between the human hand (or arm) and the manipulated object, while disregarding irrelevant background elements or distractions. Mimicking this selective attention mechanism could enhance the efficiency and effectiveness of robot learning from videos.
Drawing inspiration from this, we propose a novel approach that enables robots to learn skills by extracting and leveraging semantic representations from large-scale video collections. Our framework outlined in Fig. 1 focuses on identifying the key semantic elements relevant to skill acquisition, much like how humans naturally attend to meaningful components while learning from visual demonstrations. By concentrating on these semantic features - such as object interactions, body poses, and motion patterns - rather than processing entire scenes indiscriminately, our approach aims to make video-based skill learning more efficient and generalizable. Our key contributions are threefold:
@@
    We propose ViSA-Flow, a framework for pre-training generative policies using large-scale Video Semantic Action Flow, capturing spatio-temporal manipulator-object interactions from diverse human video demonstrations. This enables efficient knowledge transfer from Internet-scale human video data to robotic manipulation policies.
@@
    We evaluate ViSA-Flow in both simulated and real-world robotic manipulation tasks, demonstrating substantial performance improvements over SOTA baselines. Our method boosts task success rates, highlighting the effectiveness of video-driven robot skill learning.
2 Related Work
@@
Visual-Feature-Based Imitation Learning. Recent advancements [7, 8, 9, 10, 11] in visual feature-based imitation learning have significantly improved the efficiency, generalization, and robustness of learning from visual demonstrations. VIEW [12] introduces a trajectory segmentation approach that extracts condensed prior trajectories from demonstrations, allowing robots to learn manipulation tasks more efficiently. Similarly, K-VIL [13] enhances efficiency by extracting sparse, object-centric keypoints from visual demonstrations, reducing redundancy and improving learning speed. Beyond efficiency, generalization remains a critical challenge, particularly in adapting to diverse visual environments. Stem-OB [14] addresses this issue by leveraging diffusion model inversion to suppress low-level visual differences, improving robustness against variations in lighting and texture. In addition, goal-oriented approaches have been developed to improve policy learning and adaptation. Visual hindsight self-imitation learning [15] introduces hindsight goal re-labeling and prototypical goal embedding, enhancing sample efficiency in vision-based tasks.
@@
Video-Based Robot Learning. Recent advancements [16, 17, 18, 19] in robot learning have demonstrated the effectiveness of large-scale video datasets for pre-training models and improving generalization. Methods such as Time-Contrastive Networks (TCN) [20] have pioneered the extraction of temporally consistent features to align human demonstrations with robot actions. Building on this foundation, video pretraining [21] has shown that large-scale video data can be used to pretrain robust visual representations for downstream manipulation tasks. More recent works [22] have further leveraged large-scale video datasets to enhance manipulation performance. Similarly, Vid2Robot [23] presents an end-to-end framework that directly translates video demonstrations and real-time observations into robot actions, leveraging cross-attention mechanisms for improved alignment. [6] highlights the potential of leveraging partially-annotated data to enhance robot policy learning by integrating multi-modal information.
3 Method
@@
Our approach facilitates learning robot manipulation policies from limited target-domain data by leveraging knowledge distilled from large-scale source-domain (human) videos. This is achieved through the introduction and utilization of Video Semantic Action Flow (ViSA-Flow), a structured intermediate representation designed for cross-domain transfer. We first formulate the conceptual properties of ViSA-Flow and motivate its suitability for transfer learning, then detail its concrete implementation within our two-stage learning framework.
3.1 Problem Definition
@@
3.2 ViSA-Flow Representation
We propose ViSA-Flow as an intermediate representation $z_{t}\in Z_{\text{ViSA-Flow}}$ obtained by mapping an observation $o_{t}$ and context $l$ through a function $f:O\times L\rightarrow Z_{\text{ViSA-Flow}}$. The motivation is to define a representation space $Z_{\text{ViSA-Flow}}$ where the manipulation interaction relevant to the task is preserved while the domain-specific nuisance factors are mitigated, facilitating skill transfer from $O^{S}$ to $O^{T}$.
1) Semantic Entity Grounding.
Given the initial observation frame $o_{0}$ and context $l$, we utilize a pre-trained Vision-Language Model (VLM) to ground textual descriptions of the manipulator (e.g., ‘hand’, ‘gripper’) and task-relevant objects (e.g., ‘red block’) identified from $l$. A segmentation model (e.g., SAM [24]) then generates initial segmentation masks for these grounded entities, including manipulators and objects, i.e., $\{m_{M,0},m_{O_{k},0}\}$.
2) Hand-Object Interaction Tracking.
Due to the instability of semantic segmentation across sequential frames, we propose tracking the correctly segmented hand-object interaction mask over time. Specifically, we instantiate a robust point tracker (e.g., CoTracker [25]) with points densely sampled within the initial masks. The tracker estimates the 2D image trajectories $P_{t}=\{p_{j,t}\}_{j=0}^{N}$ for these points across the sequence $\{o_{t}\}_{t=0}^{T}$. These trajectories $P_{t}$ represent the extracted raw flow information, capturing the motion of key interaction points.
3) Flow-Conditioned Feature Encoding.
@@ 3.2 ViSA-Flow Representation
This mask modulates pixel intensities by an amplification factor $\alpha$ within these regions of interest, while maintaining contextual information elsewhere. The resulting perceptually-enhanced frame exhibits selective luminance amplification at interaction-critical regions. This pre-processed frame is then passed through a vision encoder $\phi$ (e.g., MAE [26]), transforming the flow-highlighted observations into our implemented ViSA-Flow representation $z_{t}$:
@@
3.3 Policy Learning through ViSA-Flow Representation
Our learning framework leverages the extracted ViSA-Flow representations $z_{t}$ within a two-stage pre-training and fine-tuning scheme, implemented using a transformer architecture, denoted $g_{\psi}$ (parameters $\psi$), inspired by prior work such as GR-1 [22].
Model Architecture.
A transformer $g_{\psi}$ is designed to process multimodal sequences for both generative prediction and policy inference shown in Fig. 2. Its input is a sequence formed by concatenating tokens representing various modalities and special learnable query tokens. Primary input modalities include language instruction embeddings $\text{Emb}(l)$ (e.g., from CLIP [27]), the sequence of recent ViSA-Flow representations $\{z_{t-h},...,z_{t}\}$ encoding flow-conditioned visual features (Sec. 3.2), the sequence of proprioceptive states $\{s_{t-h},...,s_{t}\}$ (processed via linear embeddings), and potentially tokens representing a goal state $z_{goal}$. Added to these are special query tokens: an [ACT] token for action prediction and multiple [OBS] tokens for predicting future ViSA-Flow states. Standard positional embeddings are added to this combined sequence to encode temporal order before processing by the transformer blocks. The output embeddings corresponding to the query tokens are then directed to task-specific heads; notably, the [ACT] token’s output yields the action chunk prediction $\hat{a}_{t+1:t+k}$, while the [OBS] tokens’ outputs yield predictions $\hat{z}_{t+1:t+n}$ for future states.
Stage 1: Pre-training – Learning ViSA-Flow Dynamics Prior.
Using the large-scale human video dataset $D_{v}$, we pre-train $g_{\psi}$ to model the dynamics within the ViSA-Flow space. For each sequence $v_{i}\in D_{v}$, we extract $\{z_{i,t}\}$ (Sec. 3.2). The model is trained to predict future representations $z_{t+1:t+n}$ based on past context $z_{\leq t}$ and $l$, using the [OBS] query tokens. The objective is to minimize the prediction error, typically via Mean Squared Error (MSE):
$\mathcal{L}_{\text{pretrain}}(\psi)=\mathbb{E}_{v\sim D_{v}}\left[||g_{\psi}(z_{\leq t},l)_{\text{[OBS]}}-z_{t+1:t+n}||^{2}\right].$ (3)
@@ 3.3 Policy Learning through ViSA-Flow Representation
$\mathcal{L}_{\text{finetune}}(\psi)=\mathbb{E}_{\tau\sim D_{\tau}}\left[\mathcal{L}_{\text{act}}(a_{t+1:t+k},\hat{a}_{t+1:t+k})+\lambda_{\text{fwd}}\mathcal{L}_{\text{obs}}(z_{t+1:t+n},\hat{z}_{t+1:t+n})+\lambda_{\text{prog}}\mathcal{L}_{\text{prog}}(p_{t},\hat{p}_{t})\right].$ (4)
Here, $\hat{a}_{t}=g_{\psi}(z_{\leq t},s_{\leq t},l)_{\text{[ACT]}}$ is the predicted action. $\mathcal{L}_{\text{act}}$ is the action loss (e.g., a weighted combination of Smooth L1, BCE, KL divergence terms appropriate for the action space). $\hat{z}_{t+1:t+n}=g_{\psi}(z_{\leq t},s_{\leq t},l)_{\text{[OBS]}}$ are predicted future ViSA-Flow states, and $\mathcal{L}_{\text{obs}}$ is the forward dynamics loss (MSE, identical form to Eq. 3 but on $D_{\tau}$) weighted by $\lambda_{\text{fwd}}$. $\hat{p}_{t}$ is the optional predicted progress, with $\mathcal{L}_{\text{prog}}$ being the progress loss (e.g., MSE) weighted by $\lambda_{\text{prog}}$. This stage adapts the general dynamics prior to the specific robot and learns the mapping from ViSA-Flow states (and proprioception) to robot actions, yielding the final policy parameters $\psi$.
4 Evaluation
We conduct extensive experiments in both simulated and real-world environments to systematically evaluate ViSA-Flow’s performance. Our evaluation is designed to answer the following key questions: 1) Can ViSA-Flow effectively learn and generalize across multiple tasks, particularly in challenging scenarios involving distractors, different backgrounds, and new objects? 2) Can ViSA-Flow effectively learn and generalize across diverse tasks using minimal expert demonstration data, particularly in scenarios where expert demonstration data with language annotations are scarce? 3) Do semantic actions extracted from human demonstrations benefit robot skill learning?
4.1 Simulation Experiments

### v3

@@ Abstract
I Introduction
@@
Robot imitation learning has achieved remarkable success in enabling robots to acquire complex manipulation skills, ranging from basic object manipulation[hussein2017imitation, gao2024prime] to intricate assembly procedures[chi2023diffusion]. However, the scalability of imitation learning approaches is fundamentally limited by the need for extensive, carefully curated robot datasets that are costly to collect. This has become a critical bottleneck in developing robots capable of performing diverse real-world tasks.
@@
In contrast, humans demonstrate an extraordinary ability to learn new skills by observing others. From in-person observation to videos, humans naturally focus on semantically relevant components. For instance, when learning tennis, we naturally attend to the player’s body movements, racquet handling techniques, and ball trajectories, while effectively filtering out irrelevant background information. This selective attention to meaningful elements enables efficient skill acquisition and transfer. The vast repository of publicly available videos on the internet similarly represents an untapped resource for robot learning, offering diverse demonstrations of human skills across countless domains. However, effectively leveraging this resource requires addressing several key challenges, particularly in bridging the gap between human demonstrations in unconstrained videos and robot execution in the real world.
@@
Figure 1: Humans and robots often share underlying atomic actions for similar tasks. Our framework leverages large-scale, unlabeled human videos by extracting weakly supervised semantic action flow priors (ViSA-Flow). This knowledge is distilled into a human policy and efficiently transferred to learn a corresponding robot policy.
@@
Recent research[bahl2022human, zeng2024learning, li2024grmg] has investigated how robots can acquire skills by directly observing unstructured human videos. These methods have shown promising generalization, enabling robots to adapt to new tasks beyond their training data. However, most existing approaches[wen2024anypointtrajectorymodelingpolicy, xu2024flowcrossdomainmanipulationinterface, yuan2024generalflowfoundationaffordance] rely heavily on motion flow as a conditional input for policy learning. While effective in some settings, this low-level representation often overlooks the higher-level semantic cues that humans naturally attend to when learning new skills. In real-world scenarios, when humans acquire a skill, we rarely process the entire visual scene indiscriminately. Instead, we focus selectively on the interaction between the hand (or arm) and the relevant object, while disregarding irrelevant background elements or distractions. Mimicking this selective attention mechanism could make robot learning from videos more efficient and robust.
Drawing inspiration from this observation, we propose a novel framework that learns robot skills by extracting temporally consistent semantic action flows from large-scale human manipulation videos (Fig. 1). Unlike prior works that condition policy learning on motion flow, our approach leverages semantic representations—capturing object interactions, body poses, and motion patterns—that can be consistently transferred from human demonstrations to robotic actions. By focusing on these meaningful semantic structures, our method enables more efficient and generalizable skill learning from videos. Our key contributions are threefold:
@@
    We propose ViSA-Flow, a framework for pre-training generative policies using large-scale Video Semantic Action Flow, capturing spatio-temporal manipulator-object interactions from diverse human video demonstrations. This enables efficient knowledge transfer from large-scale human video data to robotic manipulation policies and offers a new perspective on representing action sequences for cross-domain transfer.
@@
    We evaluate ViSA-Flow in both simulated and real-world robotic manipulation tasks, demonstrating substantial performance improvements over SOTA baselines, particularly in low-data regimes.
II Related Work
@@
Visual Imitation Learning. Recent advancements[tian2024predictiveinversedynamicsmodels, bu2024closedloopvisuomotorcontrolgenerative, yang2025s, mees2022matterslanguageconditionedrobotic, black2023susie] in visual feature-based imitation learning have significantly improved the efficiency, generalization of learning from visual demonstrations. VIEW[jonnavittula2025viewvisualimitationlearning] introduces a trajectory segmentation approach that extracts condensed prior trajectories from demonstrations, allowing robots to learn manipulation tasks more efficiently. Similarly, K-VIL[Gao_2023] enhances efficiency by extracting sparse, object-centric keypoints from visual demonstrations, reducing redundancy and improving learning speed. Beyond efficiency, generalization remains a critical challenge, particularly in adapting to diverse visual environments. Stem-OB[hu2024stemobgeneralizablevisualimitation] addresses this issue by leveraging diffusion model inversion to suppress low-level visual differences, improving robustness against variations in lighting and texture. In addition, goal-oriented approaches have been developed to improve policy learning and adaptation. Visual hindsight self-imitation learning[Kim_2024] introduces hindsight goal re-labeling and prototypical goal embedding, enhancing sample efficiency in vision-based tasks.
@@
Video-Based Robot Learning. Recent advancements[brohan2023rt2, reuss2024multimodaldiffusiontransformerlearning, zhou2024languageconditionedimitationlearningbase, li2024visionlanguagefoundationmodelseffective] in robot learning have demonstrated the effectiveness of large-scale video datasets for pre-training models and improving generalization. Methods such as Time-Contrastive Networks (TCN)[sermanet2018timecontrastivenetworksselfsupervisedlearning] have pioneered the extraction of temporally consistent features to align human demonstrations with robot actions. Building on this foundation, video pretraining[baker2022videopretrainingvptlearning] has shown that large-scale video data can be used to pretrain robust visual representations for downstream manipulation tasks. More recent works[wu2023unleashinglargescalevideogenerative] have further leveraged large-scale video datasets to enhance manipulation performance. Similarly, Vid2Robot [jain2024vid2robotendtoendvideoconditionedpolicy] directly translates video demonstrations into robot actions using cross-attention for alignment. [li2024grmg] highlights the potential of leveraging partially-annotated data to enhance robot policy learning by integrating multi-modal information. [lepert2503phantom] transforms human video demonstrations into robot-compatible observation-action pairs by inpainting the human arm, and overlaying a rendered robot to achieve visual domain alignment. Beyond task-specific learning, Ye et al. [ye2025video2policyscalingmanipulationtasks] explored scaling up robot learning via Internet videos, investigating how web-scale human video datasets can enhance policy learning efficiency.
Flow-Guided Imitation Learning. A growing body of work has explored representing manipulation trajectories in more flexible and generalizable forms to enhance policy learning. Wen et al.[wen2024anypointtrajectorymodelingpolicy] introduced Any-point Trajectory Modeling (ATM), which allows policies to query and generate trajectory states at arbitrary temporal points. This continuous-time representation enables efficient interpolation and improved temporal flexibility, reducing dependence on fixed-horizon action sequences. Bharadhwaj et al.[bharadhwaj2024track2actpredictingpointtracks] proposed Track2Act, which leverages point tracks extracted from Internet videos to learn generalizable manipulation skills. By grounding policies in persistent motion tracks, their method facilitates robust transfer across embodiments and visual domains without requiring paired robot demonstrations.
Another emerging line of research focuses on using flow-based representations as a domain-agnostic interface for manipulation. Xu et al.[xu2024flowcrossdomainmanipulationinterface] introduced a method that predicts dense optical flow fields as an intermediate representation, enabling the transfer of skills across domains by decoupling perception from control. Extending this idea, Yuan et al.[yuan2024generalflowfoundationaffordance] proposed general flow as a unified affordance representation for scalable robot learning, demonstrating its capacity to generalize across object categories and manipulation tasks. Flow-based interfaces effectively capture motion intent while enabling cross-domain consistency. Ren et al.[ren2025motiontracksunifiedrepresentation] introduced Motion Tracks, an agent-agnostic trajectory representation for few-shot human-to-robot transfer, enabling rapid adaptation to new tasks with minimal robot-specific data.
All these works leverage dense optical flow or motion tracks as intermediate representations, enabling skill transfer across domains. These approaches have shown strong results using non-robot videos to improve robot policies. Our method differs in the type of representation: rather than relying on flow fields, we use weak hand–object segmentation masks amplified by temporal tracking (Sec. III), aiming to capture higher-level semantic interactions.
III Method
@@
Our approach facilitates learning robot manipulation policies from limited target-domain data by leveraging knowledge distilled from large-scale source-domain (human) videos. This is achieved through the introduction and utilization of Video Semantic Action Flow (ViSA-Flow), a structured intermediate representation designed for cross-domain transfer. We first formulate the conceptual properties of ViSA-Flow and motivate its suitability for transfer learning, then detail its concrete implementation via our two-stage learning framework.
III-A Problem Definition
@@
III-B ViSA-Flow Representation
We propose ViSA-Flow as an intermediate representation $z_{t}\in Z_{\text{ViSA-Flow}}$ obtained by mapping an observation $o_{t}$ and context $l$ through a function $f:O\times L\rightarrow Z_{\text{ViSA-Flow}}$. We design $Z_{\text{ViSA-Flow}}$ to preserve task-relevant interactions while mitigating domain-specific nuisance factors, enabling cross-domain transfer.
Semantic Entity Grounding.
Given the initial observation frame $o_{0}$ and context $l$, we utilize a pre-trained Vision-Language Model (VLM) to ground textual descriptions of the manipulator (e.g., ‘hand’, ‘gripper’) and task-relevant objects (e.g., ‘red block’) identified from $l$. A segmentation model (e.g., SAM[kirillov2023segment]) then generates initial segmentation masks for these grounded entities, including manipulators and objects, i.e., $\{m_{M,0},m_{O_{k},0}\}$.
Hand-Object Interaction Tracking.
Due to the instability of semantic segmentation across sequential frames, we propose tracking the correctly segmented hand-object interaction mask over time. Specifically, we instantiate a robust point tracker (e.g., CoTracker[karaev2024cotracker3simplerbetterpoint]) with points densely sampled within the initial masks. The tracker estimates the 2D image trajectories $P_{t}=\{p_{j,t}\}_{j=0}^{J}$ for these points across the sequence $\{o_{t}\}_{t=0}^{T}$. These trajectories $P_{t}$ represent the extracted raw flow information.
Flow-Conditioned Feature Encoding.
@@ III-B ViSA-Flow Representation
This mask modulates pixel intensities by an amplification factor $\alpha$ within these regions of interest, while maintaining contextual information elsewhere. The resulting perceptually-enhanced frame exhibits selective luminance amplification at interaction-critical regions. This pre-processed frame is then passed through a pre-trained vision encoder $\phi$ (e.g., MAE[he2021maskedautoencodersscalablevision]) which is frozen during policy learning, transforming the flow-highlighted observations into our implemented ViSA-Flow representation $z_{t}$:
$z_{t}$ $=\phi\!\bigl(o_{t}\odot[1+\alpha\,M_{t}\,]\bigr).$ (2)
III-C Policy Learning through ViSA-Flow Representation
Our learning framework leverages the extracted ViSA-Flow representations $z_{t}$ within a two-stage pre-training and fine-tuning scheme, implemented using a transformer architecture, denoted $g_{\psi}$ (parameters $\psi$), inspired by prior work such as GR-1[wu2023unleashinglargescalevideogenerative].
Model Architecture.
A transformer $g_{\psi}$ is designed to process multimodal sequences for both generative prediction and policy inference shown in Fig. 2. Its input is a sequence formed by concatenating tokens representing various modalities and special learnable query tokens. Primary input modalities include language instruction embeddings $\text{Emb}(l)$ (e.g., from CLIP[radford2021learningtransferablevisualmodels]), the sequence of recent ViSA-Flow representations $\{z_{t-h},...,z_{t}\}$ encoding flow-conditioned visual features (Sec. III-B), the sequence of proprioceptive states $\{s_{t-h},...,s_{t}\}$ (processed via linear embeddings), and potentially tokens representing a goal state $z_{goal}$. Added to these are special query tokens: an [ACT] token for action prediction and multiple [OBS] tokens for predicting future ViSA-Flow states. Standard positional embeddings are added to this combined sequence to encode temporal order before processing by the transformer blocks. The output embeddings corresponding to the query tokens are then directed to task-specific heads; notably, the [ACT] token’s output yields the action chunk prediction $\hat{a}_{t+1:t+k}$, while the [OBS] tokens’ outputs yield predictions $\hat{z}_{t+1:t+n}$ for future states.
Stage 1: Pre-training – Learning ViSA-Flow Dynamics Prior.
Using the large-scale human video dataset $D_{v}$, we pre-train $g_{\psi}$ to model the dynamics within the ViSA-Flow space. For each sequence $v_{i}\in D_{v}$, we extract $\{z_{i,t}\}$ (Sec. III-B). The model is trained to predict future representations $z_{t+1:t+n}$ based on past context $z_{\leq t}$ and $l$, using the [OBS] query tokens. The objective is to minimize the prediction error, typically via Mean Squared Error (MSE):
@@ III-C Policy Learning through ViSA-Flow Representation
$\begin{split}\mathcal{L}_{\text{finetune}}(\psi)=\mathbb{E}_{\tau\sim D_{\tau}}\Bigl[\mathcal{L}_{\text{act}}\bigl(a_{t+1:t+k},\,\hat{a}_{t+1:t+k}\bigr)\\ {}+\lambda_{\text{fwd}}\,\mathcal{L}_{\text{obs}}\bigl(z_{t+1:t+n},\,\hat{z}_{t+1:t+n}\bigr)+\lambda_{\text{prog}}\,\mathcal{L}_{\text{prog}}\bigl(p_{t},\,\hat{p}_{t}\bigr)\Bigr]\end{split}$ (4)
Here, $\hat{a}_{t}=g_{\psi}(z_{\leq t},s_{\leq t},l)_{\text{[ACT]}}$ is the predicted action. $\mathcal{L}_{\text{act}}$ is the action loss combining Smooth L1 (joint regression), BCE (gripper command), and KL divergence (distribution regularization). $\hat{z}_{t+1:t+n}=g_{\psi}(z_{\leq t},s_{\leq t},l)_{\text{[OBS]}}$ are predicted future ViSA-Flow states, and $\mathcal{L}_{\text{obs}}$ is the forward dynamics loss (MSE, identical form to Eq. 3 but on $D_{\tau}$) weighted by $\lambda_{\text{fwd}}$. $\hat{p}_{t}$ is the optional predicted progress, with $\mathcal{L}_{\text{prog}}$ being the progress loss (e.g., MSE) weighted by $\lambda_{\text{prog}}$. This stage adapts the general dynamics prior to the specific robot and learns the mapping from ViSA-Flow states (and proprioception) to robot actions, yielding the final policy parameters $\psi$.
IV Evaluation
Figure 3: Datasets used for pretraining, finetuning, and evaluation. We pretrain on Something-Something-V2 with text labels and placeholders to extract semantic action flow. We finetune on 34 tasks across CALVIN environments A–C and evaluate zero-shot on environment D [mees2022calvin], where the robot completes 5 consecutive subtasks in one sequence.
We conduct extensive experiments in both simulated and real-world environments to systematically evaluate ViSA-Flow’s performance. Our evaluation is designed to answer the following key questions: 1) Can ViSA-Flow generalize across tasks with distractors, different backgrounds, and novel objects? 2) Can ViSA-Flow effectively learn and generalize across diverse tasks when expert demonstration data with language annotations are scarce? 3) Do semantic actions from human videos benefit robot skill learning?
IV-A Simulation Experiments